# High-Order Hermite Optimization: Fast and Exact Gradient Computation in Open-Loop Quantum Optimal Control using a Discrete Adjoint Approach

### v2

@@ 3.1 Hermite Runge-Kutta Methods
$\frac{d}{dt}{\mathbf{w}}(t)=A(t;\boldsymbol{\theta}){\mathbf{w}}(t)\quad 0\leq t\leq T,\quad{\mathbf{w}}(0)={\mathbf{w}}_{0}\in\mathbb{R}^{N}.$ (11)
Here the matrix $A(t;\boldsymbol{\theta})$ depends on time and a vector $\boldsymbol{\theta}$ which parameterizes the the time-dependence of the controls. For each simulation of (11) the parameter vector $\boldsymbol{\theta}$ will be fixed and for notational convenience we suppress it in this sub-section and write $A(t;\boldsymbol{\theta})=A(t)$.
@@
One way of constructing a Runge-Kutta method is to introduce $s$ nodes between $t_{n}$ and $t_{n+1}$ and then integrate the interpolant of $\frac{d{\mathbf{w}}(t)}{dt}$ exactly. This is also the approach in a Hermite Runge-Kutta method where the $s=p+q$ nodes are split into two sets of nodes containing $p$ and $q$ nodes respectively. Unlike most Runge-Kutta methods, here there are no nodes in the interior of the interval $[t_{n},t_{n+1}]$ and instead $p$ nodes coalesce at $t_{n}$ and $q$ nodes coalesce at $t_{n+1}$. Then the interpolation polynomial becomes a Hermite interpolation polynomial of degree $p+q-1$. That is
@@
where $\mathfrak{L}_{0,j}(t)$ and $\mathfrak{L}_{1,j}(t)$ are the unique degree $p+q-1$ polynomials (sometimes called generalized Lagrange polynomials, see [9]) defined through the conditions
@@
After evaluating the integrals of $\mathfrak{L}_{0,j}(t)$ and $\mathfrak{L}_{1,j}(t)$ the resulting one-step method [7] takes the form
@@ 3.2 Computing Higher Order Derivatives and Solving The Linear System of Equations
For a general non-linear ODE the computation of the derivatives ${\mathbf{w}}_{n}^{(j+1)}$ can be cumbersome but when the ODE is governed by (11) this computation can be done efficiently and recursively [7, 19] by differentiating (11) $j$ times and using the generalized Leibniz rule:
@@
so that a single timestep consists of solving a single linear system of $N$ equations:
@@
We can either form the $N\times N$ matrices $L_{n+1,q}$ and $R_{n,p}$ explicitly by explicitly forming the matrices $D_{n,j}$ and then adding them as in (17), or we compute $L_{n+1,q}{\mathbf{w}}_{n+1}$ and $R_{n,p}{\mathbf{w}}_{n}$ in a “matrix-free” fashion by applying $D_{n+1,j}$ and $D_{n,j}$ to ${\mathbf{w}}_{n+1}$ and ${\mathbf{w}}_{n}$, and adding the resulting vectors. Furthermore, by storing $D_{n+1,j}{\mathbf{w}}_{n+1}$ and $D_{n,j}{\mathbf{w}}_{n}$ after we compute them in the first $j$ terms of the summations in (15), we can compute $D_{n+1,j+1}{\mathbf{w}}_{n+1}$ and $D_{n,j+1}{\mathbf{w}}_{n}$ with only an additional $j$ matrix-vector multiplications. Therefore it costs $q(q+1)/2$ and $p(p+1)/2$ matrix-vector multiplications to compute $L_{n+1,q}{\mathbf{w}}_{n+1}$ and $R_{n+1,q}{\mathbf{w}}_{n}$, respectively, so the cost scales quadratically with the order of the method.
In this paper, we always use the matrix-free version. We store only the matrices $A_{n}^{(j)}$, which are sparse for most quantum systems, and solve (16) iteratively, using Algorithm 1 to apply $-L_{n+1,q}$ and $R_{n,p}$.
Algorithm 1 Applying $-L_{n,q}$ (or $R_{n,p})$ to ${\mathbf{w}}_{n}$.
For the system of ODEs $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}}$, compute $-L_{n,q}{\mathbf{w}}_{n}$ (or $R_{n,p}{\mathbf{w}}_{n}$), as required for the timestepping rule (16). When $p=q$, the timestepping method is A-stable and has order of accuracy $2p$.
@@
return $\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\Delta t^{j}\frac{{\mathbf{w}}_{n}^{(j)}}{j!}$triangleright To apply $R_{n,p}$ instead of $L_{n,q}$, swap $p$ and $q$ and remove the $(-1)^{j}$ factor.
@@ 3.3 Control Pulse Ansatz
$\frac{d^{i}{\mathpzc c}_{j}(t)}{d{t}^{i}}=\sum_{k=1}^{n_{j}}\theta_{j,k}\frac{d^{i}{\mathpzc c}_{j,k}(t)}{d{t}^{i}},\quad\frac{\partial}{\partial\theta_{j,k}}\frac{d^{i}{\mathpzc c}_{j}(t)}{d{t}^{i}}=\frac{d^{i}{\mathpzc c}_{j,k}(t)}{d{t}^{i}},$
and we can choose basis functions ${\mathpzc c}_{j,k}$ whose time derivatives can be easily computed.
The only requirements with respect to the Hermite timestepping method are that we may compute the value of the control pulse and its time derivatives, as well as their gradients with respect to the control parameters, and that the control pulses are sufficiently smooth to match the order of accuracy of the method being used. In principle, any control pulse ansatz may be used.
As our particular choice of control pulse ansatz, we follow the approach of [2] and use B-splines (basis splines) [8] multiplied by carrier waves, noting the importance of good frequency selection in achieving low infidelities with a small number of control parameters [6]. B-splines have several properties that make them suitable for use as a control pulse ansatz. 1) The partition of unity property ensures that the maximum value of a B-spline curve does not exceed the maximum control coefficient of the basis splines, therefore a maximum amplitude of the control pulses can be limited by placing a maximum absolute value on the control coefficients as a constraint in the optimization problem. 2) The local support property makes the pulse parametrization intuitive, since each control coefficient only affects the B-spline curve in a finite, compact region. This also makes it easy to design control pulses which begin and end with amplitude zero, which can be experimentally desirable. 3) The number of continuous derivatives of the B-spline curve increases with the order of basis splines used. Consequently, we can build arbitrarily smooth control pulses by using higher-order basis splines. 4) Finally, there are established, robust, and efficient software implementations of the evaluation of basis splines and B-spline curves [8].
@@ 3.4 Discrete Adjoint
Our goal is to rewrite (23) and choose $\lambda_{n}$ such that they cancel out the sensitivities $d{\mathbf{w}}_{n}/d\theta_{k}$, leaving us with a gradient which is much easier to compute.
@@
Given the initial value problem $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}},\quad 0\leq t\leq T,\quad{\mathbf{w}}(0)={\mathbf{w}}_{0},$ and the discrete objective functional $\mathcal{J}=\sum_{n=1}^{{N_{\rm T}}}{\mathfrak{j}}({\mathbf{w}}_{n})$, efficiently compute the gradient $\nabla\mathcal{J}=[{d\mathcal{J}}/{d{\theta_{1}}},\dots,{d\mathcal{J}}/{d{\theta_{N_{\rm C}}}}]^{T}$ exactly with respect to the Hermite Runge-Kutta method using $q$ and $p$ derivatives for the implicit/explicit parts.
Forward Evolution (Solve Forward Equations)$\triangleright$ See Algorithm 1.
@@
Adjoint Evolution (Solve Terminal Condition and Adjoint Equations)$\triangleright$ See Algorithm 4.
Compute $-\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)$ explicitly.
@@
When the state is a matrix (e.g. a gate design problem), the gradient may be computed by performing the forward evolution, adjoint evolution, and gradient accumulation for each initial condition and summing the results. The computation is almost completely independent for each initial condition, except that computing ${\partial{\mathfrak{j}}_{n}}/{\partial{\mathbf{w}}_{n}}$, where ${\mathbf{w}}_{n}$ is the state vector after $n$ timesteps of a single intial condition may require the state vector after the $n$h timestep for every initial condition. Typically, this is only the case for the final time objective ${\mathfrak{j}}_{N_{\rm T}}$ (e.g. the infidelity). Therefore, we can perfectly parallelize the forward evolution across the intial conditions, then solve the terminal condition equation (25), then perform the rest of the adjoint evolution (26) and gradient accumulation (24) in parallel.
@@
When the time-dependence of $A(t)$ has the structure (18), the gradient accumulation (24) can be rewritten so that the number of matrix-vector multiplications required is independent of the number of control parameters ${N_{\rm C}}$.
@@
Compute $\nabla_{\boldsymbol{\theta}}\left\langle L_{n+1,q}{\mathbf{w}}_{n+1}+R_{n,p}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$, in order to perform the gradient accumulation (24) efficiently when $A$ takes the usual form: $A(t;\boldsymbol{\theta})=A_{0}+c_{1}(t;\boldsymbol{\theta})A_{1}+\dots+c_{N_{\rm C}}(t;\boldsymbol{\theta})A_{N_{\rm C}}$. Assumes $p=q$, so that the recursive procedures for computing $\nabla_{\boldsymbol{\theta}}\left\langle L_{n+1,q}{\mathbf{w}}_{n+1},\boldsymbol{\lambda}_{n+1}\right\rangle$ and $\nabla_{\boldsymbol{\theta}}\left\langle R_{n,q}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ have the same depth.
@@
return $\sum_{j=1}^{q}\textsc{$\nabla$D}(\boldsymbol{\lambda}_{n+1},\ j,\ (-1)^{j+1}c_{j}^{qp}\Delta t^{j},\ c_{j}^{pq}\Delta t^{j})$
@@
Whereas the computation of $D_{n,j+1}\mathbf{x}$ can be done efficiently by storing and reusing $D_{n,i}\mathbf{x}$ for $i=1,\dots,j$, in the computation of $(D_{n,j+1})^{T}\mathbf{x}$ the vector being multiplied by each $(D_{n,i})^{T}$ changes with the values of $i$ and $j$. The vectors $A_{n}^{(i)}\mathbf{x}$, $i=0,\dots,j$ may be computed once and and reused. When this is done, the number of matrix-vector mutiplications needed to multiply a vector by $(D_{n,j+1})^{T}$ is
@@
A comparison of the number of matrix-vector multiplications required in the application of $L_{n,q}$ or $R_{n,p}$ (as in the forward evolution) and in the application of $(L_{n,q})^{T}$ or $(R_{n,p})^{T}$ (as in the adjoint evolution) when $q=p$ is given in Table 1. The number of matrix-vector multiplications required in the adjoint evolution scales exponentially with the order of the method, limiting the efficiency of the extremely high order methods. In the rest of this work, we consider method orders twelve and under, for which the number of matrix-vector multiplications required in the adjoint evolution does not exceed three times the number required in the forward evolution.
@@
Algorithm 4 Applying $-(L_{n,q})^{T}$ (or $(R_{n,p})^{T}$) to $\boldsymbol{\lambda}_{n}$.
For the system of ODEs $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}}$, given $\boldsymbol{\lambda}_{n}$, the adjoint state corresponding to the $n$-th timestep, compute $-(L_{n,q})^{T}\boldsymbol{\lambda}_{n}$ (or $(R_{n,p})^{T}\boldsymbol{\lambda}_{n}$), as required when solving the teriminal condition (25) or adjoint equations (26).
Precompute $\ \frac{1}{(j-1)!}(A_{n}^{(j-1)})^{T}\boldsymbol{\lambda}_{n}$ for $j=1,\dots,q$.
return $\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\Delta t^{j}\textsc{DT}(\boldsymbol{\lambda}_{n},j)$triangleright To apply $R_{n,p}$ instead of $L_{n,q}$, swap $p$ and $q$ and remove the $(-1)^{j}$ factor.
@@
     $\triangleright$ (When DT is not called recursively, then $\mathbf{x}=\boldsymbol{\lambda}_{n}$ and $\frac{1}{(j-i)!}(A_{n}^{(j-i)})^{T}\mathbf{x}$ is precomputed.)
@@
Because the terminal condition (25) and adjoint equations (26) require computing $({\partial{\mathfrak{j}}_{n}}/{\partial{\mathbf{w}}_{n}})^{T}$ for $n={N_{\rm T}},\dots,1$, it is necessary to either store the entire state history ${\mathbf{w}}_{0},\dots,{\mathbf{w}}_{N_{\rm T}}$ when solving the forward equations, or to solve Schrödinger’s equation backward in time while solving the adjoint equations, so that we only every store a single state vector or adjoint state vector at each timestep. We refer to this as the “memory-lean” version. In this case, the gradient accumulation (24) is also performed alongside the adjoint evolution.
There are two main drawbacks to the memory-lean version. 1) We must solve an additional system of ODEs, increasing the cost of our method by roughly 50%. 2) Because the Hermite Runge-Kutta methods we use are non-reversible, the memory-lean version computed gradient will no longer be exact with respect to the discretized forward equations. The latter drawback can be especially detrimental to the performance of quasi-Newton optimizers. As a compromise, a checkpointing approach [15] could be used to store the state vector at only a subset of of all timesteps and still computing an accurate gradient with respect to the discretized forward equations.

### v3

@@ 3.1 Hermite Runge-Kutta Methods
$\frac{d}{dt}{\mathbf{w}}(t)=A(t;\boldsymbol{\theta}){\mathbf{w}}(t)\quad 0\leq t\leq T,\quad{\mathbf{w}}(0)={\mathbf{w}}_{0}\in\mathbb{R}^{2N}.$ (11)
Here the matrix $A(t;\boldsymbol{\theta})$ depends on time and a vector $\boldsymbol{\theta}$ which parameterizes the time-dependence of the controls. For each simulation of (11) the parameter vector $\boldsymbol{\theta}$ will be fixed and for notational convenience we suppress it in this sub-section and write $A(t;\boldsymbol{\theta})=A(t)$.
@@
One way of constructing a Runge-Kutta method is to introduce $s$ nodes between $t_{n}$ and $t_{n+1}$ and then integrate the interpolant of $\frac{d{\mathbf{w}}(t)}{dt}$ exactly. This is also the approach in a Hermite Runge-Kutta method where the $s=p+q$ nodes are split into two sets of nodes containing $p$ and $q$ nodes respectively. Unlike most Runge-Kutta methods, here there are no nodes in the interior of the interval $t_{n},t_{n+1}]$ and instead $p$ nodes coalesce at $t_{n}$ and $q$ nodes coalesce at $t_{n+1}$. Then the interpolation polynomial becomes a Hermite interpolation polynomial of degree p+q1p+q-1. That is
@@
where $\mathfrak{L}_{0,j}(t)$ and $\mathfrak{L}_{1,j}(t)$ are the unique degree p+q1p+q-1 polynomials (sometimes called generalized Lagrange polynomials, see [11]) defined through the conditions
@@
Here, $\delta_{ij}$ denotes the Kronecker delta, which equals $1$ if $i=j$ and 0 otherwise. After evaluating the integrals of $\mathfrak{L}_{0,j}(t)$ and $\mathfrak{L}_{1,j}(t)$ the resulting one-step method [9] takes the form
@@ 3.2 Computing Higher Order Derivatives and Solving The Linear System of Equations
For a general non-linear ODE, computing the derivatives ${\mathbf{w}}_{n}^{(j+1)}$ can be cumbersome. But when the ODE is governed by (11), this computation can be done efficiently and recursively [9, 23] by differentiating (11) $j$ times and using the generalized Leibniz rule:
@@
so that a single timestep consists of solving a single linear system of $2N$ equations:
@@
We can either form the $2N\times 2N$ matrices $L_{n+1,q}$ and $R_{n,p}$ explicitly by explicitly forming the matrices $D_{n,j}$ and then adding them as in (17), or we can compute $L_{n+1,q}{\mathbf{w}}_{n+1}$ and $R_{n,p}{\mathbf{w}}_{n}$ in a “matrix-free” fashion by applying $D_{n+1,j}$ and $D_{n,j}$ to ${\mathbf{w}}_{n+1}$ and ${\mathbf{w}}_{n}$, and adding the resulting vectors. Furthermore, by storing $D_{n+1,j}{\mathbf{w}}_{n+1}$ and $D_{n,j}{\mathbf{w}}_{n}$ after we compute them in the first $j$ terms of the summations in (15), we can compute $D_{n+1,j+1}{\mathbf{w}}_{n+1}$ and $D_{n,j+1}{\mathbf{w}}_{n}$ with only an additional $j$ matrix-vector multiplications. Therefore it costs $q(q+1)/2$ and $p(p+1)/2$ matrix-vector multiplications to compute $L_{n+1,q}{\mathbf{w}}_{n+1}$ and $R_{n+1,q}{\mathbf{w}}_{n}$, respectively, so the cost scales quadratically with the order of the method.
In this paper, we always use the matrix-free version. We store only the matrices $A_{n}^{(j)}$, which are sparse for most quantum systems, and solve (16) iteratively, using Algorithm 1 to apply $L_{n+1,q}$ and $R_{n,p}$.
Algorithm 1 Applying $L_{n,q}$ (or $R_{n,p})$ to ${\mathbf{w}}_{n}$.
For the system of ODEs $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}}$, compute $L_{n,q}{\mathbf{w}}_{n}$ (or $R_{n,p}{\mathbf{w}}_{n}$), as required for the timestepping rule (16). When $p=q$, the timestepping method is A-stable and has order of accuracy $2p$.
@@
return $\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\Delta t^{j}\frac{{\mathbf{w}}_{n}^{(j)}}{j!}$ $\triangleright$ To apply $R_{n,p}$ instead of $L_{n,q}$, swap $p$ and $q$ and remove the $(-1)^{j}$ factor.
@@ 3.3 Control Pulse Ansatz
$\frac{d^{i}{\mathpzc c}_{j}(t)}{d{t}^{i}}=\sum_{k=1}^{n_{j}}\theta_{j,k}\frac{d^{i}{\mathpzc c}_{j,k}(t)}{d{t}^{i}}.\quad\frac{\partial}{\partial\theta_{j,k}}\frac{d^{i}{\mathpzc c}_{j}(t)}{d{t}^{i}}=\frac{d^{i}{\mathpzc c}_{j,k}(t)}{d{t}^{i}},$
In addition, we can choose basis functions ${\mathpzc c}_{j,k}$ whose time derivatives can be easily computed.
The only requirements with respect to the Hermite timestepping method are that we may compute the value of the control pulses and their time derivatives, as well as their gradients with respect to the control parameters, and that the control pulses are sufficiently smooth to match the order of accuracy of the method being used. In principle, any control pulse ansatz may be used.
As our particular choice of control pulse ansatz, we follow the approach of [2] and use B-splines (basis splines) [10] multiplied by carrier waves, noting the importance of good frequency selection in achieving low infidelities with a small number of control parameters [7]. B-splines have several properties that make them suitable for use as a control pulse ansatz. 1) The partition of unity property ensures that the maximum value of a B-spline curve does not exceed the maximum control coefficient of the basis splines. Therefore, a maximum amplitude of the control pulses can be limited by placing a maximum absolute value on the control coefficients as a constraint in the optimization problem. 2) The local support property makes the pulse parametrization intuitive, since each control coefficient only affects the B-spline curve within a finite, compact region. This also makes it easy to design control pulses which begin and end with amplitude zero, which can be experimentally desirable. 3) The number of continuous derivatives of the B-spline curve increases with the order of basis splines used. Consequently, we can build arbitrarily smooth control pulses by using higher-order basis splines. 4) Finally, there are established, robust, and efficient software implementations of the evaluation of basis splines and B-spline curves [10].
@@ 3.4 Discrete Adjoint
Our goal is to rewrite (23) and choose $\lambda_{n}$ such that they cancel out the sensitivities $d{\mathbf{w}}_{n}/d\theta_{k}$, leaving us with a gradient is much easier to compute.
@@
Given the initial value problem $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}},\quad 0\leq t\leq T,\quad{\mathbf{w}}(0)={\mathbf{w}}_{0},$ and the discrete objective functional $\mathcal{J}=\sum_{n=1}^{{N_{\rm T}}}{\mathfrak{j}}({\mathbf{w}}_{n})$, efficiently compute the gradient $\nabla\mathcal{J}=[{d\mathcal{J}}/{d{\theta_{1}}},\dots,{d\mathcal{J}}/{d{\theta_{{{N_{\rm P}}}}}}]^{T}$ exactly with respect to the Hermite Runge-Kutta method using $q$ and $p$ derivatives for the implicit/explicit parts.
[…]
@@
[…]
Compute $\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)$ explicitly.
@@
When the state is a matrix (e.g. a gate design problem), the gradient may be computed by performing the forward evolution, adjoint evolution, and gradient accumulation for each initial condition and summing the results. The computation is almost completely independent for each initial condition, except that computing ${\partial{\mathfrak{j}}_{n}}/{\partial{\mathbf{w}}_{n}}$, where ${\mathbf{w}}_{n}$ is the state vector after $n$ timesteps of a single initial condition may require the state vector after the $n$-th timestep for every initial condition. Typically, this is only the case for the final time objective ${\mathfrak{j}}_{N_{\rm T}}$ (e.g. the infidelity). Therefore, we can perfectly parallelize the forward evolution across the initial conditions, then solve the terminal condition equation (25), then perform the rest of the adjoint evolution (26) and gradient accumulation (24) in parallel.
@@
When the time-dependence of $A(t)$ has the structure (18), the gradient accumulation (24) can be rewritten so that the number of matrix-vector multiplications required is independent of the number of control parameters $N_{\rm P}$}}.
@@
Compute $\nabla_{\boldsymbol{\theta}}\left\langle L_{n+1,q}{\mathbf{w}}_{n+1}+R_{n,p}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$, in order to perform the gradient accumulation (24) efficiently when AA takes the usual form: $A(t;\boldsymbol{\theta})=A_{0}+c_{1}(t;\boldsymbol{\theta})A_{1}+\dots+c_{N_{\rm C}}(t;\boldsymbol{\theta})A_{N_{\rm C}}$. Assumes $p=q$, so that the recursive procedures for computing $\nabla_{\boldsymbol{\theta}}\left\langle L_{n+1,q}{\mathbf{w}}_{n+1},\boldsymbol{\lambda}_{n+1}\right\rangle$ and $\nabla_{\boldsymbol{\theta}}\left\langle R_{n,q}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ have the same depth.
@@
[…]
@@
Whereas the computation of $D_{n,j+1}\mathbf{x}$ can be done efficiently by storing and reusing $D_{n,i}\mathbf{x}$ for $i=1,\dots,j$, in the computation of $(D_{n,j+1})^{T}\mathbf{x}$ the vector being multiplied by each $(D_{n,i})^{T}$ changes with the values of $i$ and $j$. The vectors $A_{n}^{(i)}\mathbf{x}$, $i=0,\dots,j$ may be computed once and and reused. When this is done, the number of matrix-vector multiplications needed to multiply a vector by $(D_{n,j+1})^{T}$ is
@@
A comparison of the number of matrix-vector multiplications required in the application of $L_{n,q}$ or $R_{n,p}$ (as in the forward evolution) and in the application of $(L_{n,q})^{T}$ or $(R_{n,p})^{T}$ (as in the adjoint evolution) when $q=p$ is given in Table 2. The number of matrix-vector multiplications required in the adjoint evolution scales exponentially with the order of the method, limiting the efficiency of the extremely high-order methods. In the rest of this work, we consider method orders twelve and under, for which the number of matrix-vector multiplications required in the adjoint evolution does not exceed three times the number required in the forward evolution.
@@
Algorithm 4 Applying $(L_{n,q})^{T}$ (or $(R_{n,p})^{T}$) to $\boldsymbol{\lambda}_{n}$.
For the system of ODEs $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}}$, given $\boldsymbol{\lambda}_{n}$, the adjoint state corresponding to the $n$-th timestep, compute $(L_{n,q})^{T}\boldsymbol{\lambda}_{n}$ (or $(R_{n,p})^{T}\boldsymbol{\lambda}_{n}$), as required when solving the teriminal condition (25) or adjoint equations (26).
[…]
return $\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\Delta t^{j}\textsc{DT}(\boldsymbol{\lambda}_{n},j)$ $\triangleright$ To apply $R_{n,p}$ instead of $L_{n,q}$, swap $p$ and $q$ and remove the $(-1)^{j}$ factor.
@@
[…]
@@
Because the terminal condition (25) and adjoint equations (26) require computing $({\partial{\mathfrak{j}}_{n}}/{\partial{\mathbf{w}}_{n}})^{T}$ for $n={N_{\rm T}},\dots,1$, it is necessary to either store the entire state history ${\mathbf{w}}_{0},\dots,{\mathbf{w}}_{N_{\rm T}}$ when solving the forward equations, or to solve Schrödinger’s equation backward in time while solving the adjoint equations, so that we only every store a single state vector or adjoint state vector at each timestep. We refer to the latter approach as the “memory-lean” version. In this case, the gradient accumulation (24) is also performed alongside the adjoint evolution.
There are two main drawbacks to the memory-lean version. 1) We must solve an additional system of ODEs, increasing the cost of our method by roughly 50%. 2) Because the Hermite Runge-Kutta methods we use are non-reversible, the memory-lean version computed gradient will no longer be exact with respect to the discretized forward equations. The latter drawback can be especially detrimental, depending on the magnitude of the error introduced by the memory-lean version, since the gradient becomes smaller as the objective function is minimized. As the length of the gradient decreases, eventually the approximation error dominates the computed value of the gradient. When using the GRAPE approach with an inexact gradient formula, de Fouquieres et al. [13] observed that the convergence of the optimization slows down dramatically compared to the convergence when using an exact gradient as the optimization reaches high fidelities (the threshold at which the slowdown occurs increases with the accuracy of the gradient), effectively imposing a maximum fidelity achievable by the optimization algorithm.
As a compromise, a checkpointing approach [19] could be used to store the state vector at only a fraction of all the timesteps but still compute an exact gradient. In order to recover the states between the checkpoints, the forward equations must be solved a second time while solving the adjoint equations. Similar to the memory-lean version, this increases the cost of the method by roughly 50%. In the numerical examples that follow, we always store the entire state history.